OpenStudy (anonymous):

Find the mean of the following data: 7, -6, 8, -10, -8, 4, 5

OpenStudy (katrinakaif):

Add them all up, and divide by the number of terms there are in the set..so in this case (sum of all numbers) / 7

OpenStudy (anonymous):

answer is 0

OpenStudy (anonymous):

mean = (7-6+8-10-8+4+5)/7 =0

OpenStudy (anonymous):

thanks!

OpenStudy (katrinakaif):

Anytime =)

OpenStudy (anonymous):

add all the numbers, you get an answers COUNT how many numbers you have added and divded the number you got that you added EX: 2+2=4 4divided by 2 =2. The mean is almost known as the average